A Comparison of Implicitly Parallel Multithreaded and Data Parallel Implementations of an Ocean Model

Computation Structures Group Memo 364 September 24, 1997

Andrew Shaw Arvind Kyoo-Chan Cho Christopher Hill R. Paul Johnson John Marshall

This report describes research done at the Laboratory for Computer Science and the Department of Earth, Atmospheric, and Planetary Sciences of the Massachusetts Institute of Technology. Funding for the Laboratory for Computer Science is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Of�ce of Naval Research contract N00014-92-J-1310.

1